OpenStudy (anonymous):

Change 2x + 5y = -30 to slope-intercept form and state the slope and y-intercept. You must show all work.

OpenStudy (smurfy14):

put it in y=mx+b form. 5y=-2x-30 Now divide by 5 Answer: y=-2/5x-6